OpenStudy (anonymous):

144% of what number is 126?

OpenStudy (theedwardsfamily):

what do you think?

OpenStudy (anonymous):

126/144?

OpenStudy (stamp):

144% * n = 126 144/100 * n = 126 1.44 * n = 126 n = 126/1.44

OpenStudy (theedwardsfamily):

exactly^

OpenStudy (anonymous):

oh okay thank you everyone

OpenStudy (theedwardsfamily):

126/144=87.5 OR 87.5%

OpenStudy (anonymous):

i got 87.5

OpenStudy (theedwardsfamily):

yes